# Kleene Algebrae, Kleene Modules, and Morita Equivalence

### v1

@@ 2.1 The Morita Category
Semirings $R$ and $S$ are Morita-equivalent if and only if for some dimension $n$ and for some full idempotent $e$ in the semiring $M$ of $n\times n$ matrices over the ring (without loss of generality) $R$, $S\cong eMe$.
Certainly Morita-equivalent Kleene algebrae are Morita equivalent as semirings, and so any Kleene algebra $S$ Morita-equivalent to the algebra $K$ is, up to isomorphism, of the form $eMe$ for some finite-dimensional matrix ring $R$ over $K$ and some full idempotent $e\in M$. It is natural to ask whether Kleene algebrae whose semiring reducts are Morita equivalent must be Morita equivalent as Kleene algebrae. That is to say, if there is a semimodule Morita isomorphism between Kleene algebrae $K$ and $S$, must there also be a Kleene algebra Morita isomorphism?
Proposition 7.
Let $K$ and $S$ be Kleene algebrae. If the semiring reducts of $K$ and $S$ are Morita equivalent, then $K$ and $S$ are Morita equivalent as Kleene algebrae.
@@
This appears to be hard to do directly, though maybe I just wasn’t thinking in the right direction. In any case, it is straightforward given the characterization of Morita-equivalence for semirings given above. Without loss of generality, fix a dimension $n$ such that, letting $M$ be the Kleene algebra of $n\times n$ matrices over $K$, there is a full idempotent of $M$ such that $S\cong eMe$. Then $K^{n\circ}e\otimes eK^{n}$ is a submodule of $K^{n\circ}\otimes K^{n}$, so because the idempotent $e$ is full, in fact $K^{n\circ}e\otimes eK^{n}\cong K$. Conversely,
@@
Therefore the Kleene modules $K^{n\circ}e$ and $eK^{n}$ witness that $K$ and $M$ are Morita equivalent. ∎
Acknowledgement. Thank you to Dexter Kozen for many stimulating discussions of this material, and to Goktug for helping to finalize a project direction.

### v2

@@ 4 The Morita Skeleton of Semirings and of Kleene Algebras
Semirings $R$ and $S$ are Morita-equivalent if and only if one of the following occurs for some dimension $n$:
  1. 1.
    for some full idempotent $e$ in the semiring $M$ of $n\times n$ matrices over the semiring $R$, $S\cong eMe$, or else
  2. 2.
    for some full idempotent $e^{\prime}$ in the semiring $M^{\prime}$ of $n\times n$ matrices over the semiring $S$, $R\cong e^{\prime}M^{\prime}e^{\prime}$.
Certainly Morita-equivalent Kleene algebras are Morita equivalent as semirings, and so any Kleene algebra $S$ Morita-equivalent to the algebra $K$ is, up to isomorphism, of the form $eMe$ for some finite-dimensional matrix ring $R$ over $K$ and some full idempotent $e\in M$. It is natural to ask whether Kleene algebras whose semiring reducts are Morita equivalent must be Morita equivalent as Kleene algebras. That is to say, if there is a semimodule Morita isomorphism between Kleene algebras $K$ and $S$, must there also be a Kleene algebra Morita isomorphism?
Theorem 22 (Morita rigidity of Kleene algebras over semirings).
Let $K$ and $S$ be a Kleene algebras. If the semiring reducts of $K$ and $S$ are Morita-equivalent via semiring bimodules, then in fact $K$ and $S$ are Morita-equivalent via Kleene bimodules..
The statement of this theorem is exactly what we mean when stating that the category of Kleene algebras is Morita-rigid with respect to the category of semirings.
@@
Using proposition 21, we may assume without loss of generality that for some dimension $n$, $S$ is isomorphic to $eMe$ for some full idempotent $e$ of the Kleene algebra $M$ of matrices over $S$. The asterate of $M$ is the standard matrix asterate for Kleene algebras; see [11] for the definition and basic properties. Note that $K^{n\circ}e\otimes eK^{n}$ is a submodule of $K^{n\circ}\otimes K^{n}$, so because the idempotent $e$ is full, in fact $K^{n\circ}e\otimes eK^{n}\cong K$. Conversely,
@@
Therefore the Kleene modules $K^{n\circ}e$ and $eK^{n}$ witness that $K$ and $S$ are Morita-equivalent. ∎
5 Conclusion and further directions
At the time of this writing Morita equivalence is a new theoretical tool being introduced into the study of Kleene algebras and related structures, but it does have some intriguing potential for applications to the theory of programming language semantics. One observation is that it is possible to freely change scalars from one Kleene-type algebra $K$ to a Morita equivalent one $S$, because if bimodules ${{}_{K}}M_{S}$ and ${{}_{S}}N_{K}$ witness the Morita equivalence of $K$ and $S$, then for any left module ${{}_{K}}P$ we can reversibly pass to the $S$-module ${{}_{S}}P={{}_{S}}N_{K}\otimes{{}_{K}}P$, the reversibility being witnessed by the fact that
${{}_{K}}M_{S}\otimes{{}_{S}}P\cong({{}_{K}}M_{S}\otimes{{}_{S}}N_{K})\otimes{{}_{K}}P\cong{{}_{K}}K_{K}\otimes{{}_{K}}P\cong{{}_{K}}P.$
Changes of scalars for left modules and bimodules are handled similarly.
It is also of interest to investigate what additional structure can be added to a Kleene algebra (such as order-meets, to form a Kleene lattice, or order-residuals, to form an action algebra; see [12]) while preserving Morita rigidity in the sense that Morita equivalence of semiring reducts implies Morita equivalence of the full algebras (via bimodules which in some appropriate sense preserve the additional structure). This question was suggested to the author by Jose Meseguer at Logic and Applications 2025 in Dubrovník, Croatia The author has investigated this problem but removed results stated in an earlier version due to subtleties concerning homomorphism modules and tensor products pointed out by a diligent, anonymous referee.
This work is just the beginning of what could be a large research program. One exciting direction is to develop homology and cohomology theory for Kleene algebras and investigate what implications the presence of nontrivial (co)homology classes have for computations represented in a given algebra.